### v1

Abstract
We investigate the semigroup of integer points inside a convex cone. We extend classical results in integer linear programming to integer conic programming. We show that the semigroup associated with nonpolyhedral cones can sometimes have a notion of finite generating set. We show this is true for the cone of positive semidefinite matrices (PSD) and the second-order cone (SOC). Both cones have a finite generating set of integer points, similar in spirit to Hilbert bases, under the action of a finitely generated group. We also extend notions of total dual integrality, Gomory-Chvátal closure, and Carathéodory rank to integer points in arbitrary cones.
Keywords:
Integer Points Convex Cones Semigroups Hilbert bases Conic Programming Positive Semidefinite Cone Second-Order Cone
1 Introduction
A semigroup $S$ is a subset of $\mathbb{Z}^{N}$ that contains $\mathbf{0}$ and is closed under addition. Given a convex cone $C\subseteq\mathbb{R}^{N}$, the integer points $S_{C}\mathrel{\mathop{\ordinarycolon}}=C\cap\mathbb{Z}^{N}$ form a semigroup which we will call the conical semigroup of $C$. In particular, given any compact convex body $K\subseteq\mathbb{R}^{n}$, the integer points $\text{cone}(K\times\{1\})\cap\mathbb{Z}^{n+1}$ form a conical semigroup. Conical semigroups appear not just in optimization [1, 6], but also in algebra and number theory [2, 3]. Given a convex cone $C\subseteq\mathbb{R}^{N}$ for $N\geq 1$, we say a subset $B\subseteq S_{C}$ is an integral generating set of $S_{C}$ if for any $s\in S_{C}$ there exist $b_{1},\dots,b_{m}\in B$ and $c_{1},\dots,c_{m}\in\mathbb{Z}_{\geq 0}$ such that $s=\sum_{i=1}^{m}c_{i}b_{i}$, for some $m\geq 1$. Furthermore, we call $B$ a conical Hilbert basis if $B$ is an inclusion-minimal integral generating set.
When the defining cone $C$ is rational, polyhedral and pointed, there is abundant literature on the topic. It is well-known that we have a unique finite Hilbert basis in this case [11, 20]. Historically, Hilbert bases have been fundamental in the theory and algorithms of combinatorial optimization. For example, determining if a rational system $A\mathbf{x}\leq\mathbf{b}$ is totally dual integral (TDI) is equivalent to checking if, for every face $F$ of the polyhedron $P\mathrel{\mathop{\ordinarycolon}}=\{\mathbf{x}\mathrel{\mathop{\ordinarycolon}}A\mathbf{x}\leq\mathbf{b}\}$, the rows of $A$ which are active in a face $F$ form a Hilbert basis for cone($F$)[20].
It is natural to ask, what properties transfer from rational polyhedral cones to arbitrary convex cones? For instance, do we preserve any notion of finiteness in generating sets for semigroups when we relax the polyhedral condition and instead consider general conical semigroups? Are there Hilbert bases for general cones? This paper discusses finite generation for conical semigroups and extends the polyhedral cone theory of Hilbert bases to nonpolyhedral convex cones. Our main results will pertain to the semigroups arising from the cone of positive semidefinite matrices and the second-order cone. Both cones play a key role in modern optimization [4, 5]. We also discuss some applications of our nonpolyhedral point of view.
In what follows, we denote $\mathrm{GL}(N,\mathbb{Z})\mathrel{\mathop{\ordinarycolon}}=\{U\in\mathbb{Z}^{N\times N}\mathrel{\mathop{\ordinarycolon}}~{}|\det(U)|=1\}$. Here is a new notion of finite generation for conical semigroups.
Definition 1
Given a conical semigroup $S_{C}\subset\mathbb{Z}^{N}$, we call it $(R,G)$-finitely generated if there is a finite subset $R\subseteq S_{C}$ and a finitely generated subgroup $G\subseteq\mathrm{GL}(N,\mathbb{Z})$ acting on $C$ linearly such that
  1. 1.
    both the cone $C$ and the semigroup $S_{C}$ are invariant under the group action, i.e., $G\cdot C=C$ and $G\cdot S_{C}=S_{C}$, and
  2. 2.
    every element $s\in S_{C}$ can be represented as
    $s=\displaystyle\sum_{i\in K}\lambda_{i}g_{i}\cdot r_{i}$
    for $r_{i}\in R$, $g_{i}\in G$, and $\lambda_{i}\in\mathbb{Z}_{\geq 0}$, and where $K$ is a finite index set.
Note that when $C$ is a (pointed) rational polyhedral cone, then the conical semigroup $S_{C}=C\cap\mathbb{Z}^{N}$ is $(R,G)$-finitely generated by $R$, its Hilbert basis, and $G$, the trivial group $\{I_{N}\}$. Similarly, note that if $S_{C}$ is an $(R,G)$-finitely generated semigroup, then $\cup_{r\in R}G\cdot r$ is an integral generating set of $S_{C}$, which is a superset of a conical Hilbert basis. We call $R$ the set of roots of $S_{C}$, and $\cup_{r\in R}G\cdot r$ the set of generators for $S_{C}$.
While a nonpolyhedral cone cannot be finitely generated in the usual sense, using a possibly infinite (finitely generated) group $G$ allows us to extend our understanding beyond the polyhedral case. Because the possibly infinite generators for $S_{C}$ can be obtained by group action $G$ on a finite set $R$ and $G$ is finitely generated, this allows for the possibility of algorithmic methods. The well-known Krein-Milman theorem states that any point in a closed pointed cone $C$ can be generated by extreme rays, denoted by $\operatorname{ext}(C)$ [4]. When we restrict to the conical semigroup $S_{C}$ and nonnegative integer combinations, the primitive integer points on the extreme rays of $C$ must be contained in the set of generators of $S_{C}$, where an integer point $x=(x_{1},\dots,x_{N})\in\mathbb{Z}^{N}$ is primitive if $\mathrm{gcd}(x_{1},\dots,x_{N})=1$. We call the integer points of $S_{C}$ on the extreme rays of $C$ extreme points, denoted by $\operatorname{ext}(S_{C})\mathrel{\mathop{\ordinarycolon}}=\{y\in S_{C}\mathrel{\mathop{\ordinarycolon}}y\in\operatorname{ext}(C)\cap\mathbb{Z}^{N}\}$. However, as in the polyhedral case, the generators will often include extra nonextreme boundary points or even interior points. We provide the following definition of sporadic points that cannot have an extreme point subtracted from them and still remain within the cone.
Definition 2
We say a point $x\in S_{C}=C\cap\mathbb{Z}^{N}$ is sporadic if there does not exist $y\in\operatorname{ext}(S_{C})$ such that $x-y\in S_{C}$.
If $x\in S_{C}$ is sporadic, then $x$ cannot be written as an integer conical combination of extreme points (even though it can be written as a real combination of them). From the definition of sporadic points, we know that all points $x\in S_{C}$ can be written as an integer conical combination of primitive extreme points and one sporadic point. To show that a semigroup is $(R,G)$-finitely generated, it is sufficient to show that the set of primitive extreme points and sporadic points are finite or can be obtained from a finitely generated group $G$ that acts on a finite set of roots, $R$.
The two convex cones of interest in this work are positive semidefinite cone (PSD) and second-order cone (SOC). In Sections 2 and 3 of this paper, we will present the following two main results pertaining to integer points in the PSD cone $\mathcal{S}^{n}_{+}(\mathbb{Z})$, and those in the SOC $\mathrm{SOC}(n)\cap\mathbb{Z}^{n}$.
Theorem 1.1
The conical semigroup of the cone of $n\times n$ positive semidefinite matrices, $\mathcal{S}^{n}_{+}(\mathbb{Z})$, is $(R,G)$-finitely generated by $G\cong\mathrm{GL}(n,\mathbb{Z})$ where $G$ acts on $X\in\mathcal{S}^{n}_{+}(\mathbb{Z})$ by $X\mapsto UXU^{\mathsf{T}}$ for each $U\in\mathrm{GL}(n,\mathbb{Z})$, and by $R$, the union of any single rank-one matrix and a finite subset of the sporadic points. Moreover,
  1. 1.
    If $n\leq 5$, then there are no sporadic points. Thus, $R=\{\mathbf{e}_{1}\mathbf{e}_{1}^{\mathsf{T}}\}$, where $\mathbf{e}_{1}$ is the first unit vector.
  2. 2.
    If $n=6$, then $R=\{\mathbf{e}_{1}\mathbf{e}_{1}^{\mathsf{T}},M\}$, where $M$ is a single sporadic point defined in Section 2 Proposition 5.
Theorem 1.2
For dimension $3\leq n\leq 10$, the conical semigroup $\mathrm{SOC}(n)\cap\mathbb{Z}^{n}$ is $(R,G)$-finitely generated with $G$ and $R$ defined in Section 3.
We say that two matrices $X_{1},X_{2}$ are unimodularly equivalent if $X_{2}=U\cdot X_{1}=UX_{1}U^{\mathsf{T}}$ for some $U\in\mathrm{GL}(n,\mathbb{Z})$. It is easy to see that it defines an equivalence relation for all integer PSD matrices. Note that the equivalence class of $\mathbf{e}_{1}\mathbf{e}_{1}^{\mathsf{T}}$ are all rank-1 integer matrix $\mathbf{x}\mathbf{x}^{\mathsf{T}}$ for some primitive integer vector $\mathbf{x}\in\mathbb{Z}^{n}$. An interpretation of Theorem 1.1 is that for dimension $n\leq 5$, every integer PSD matrix can be represented as the sum of rank-$1$ matrices $\mathbf{x}\mathbf{x}^{\mathsf{T}}$ for some primitive integer vector $\mathbf{x}\in\mathbb{Z}^{n}$. However, the same result fails for dimension $n=6$. In this case, we will have that every integer PSD matrix can be represented as the sum of rank-$1$ matrices and one sporadic matrix $Y$, which is unimodularly equivalent to $M$ (this matrix was first found by [19]). In general, every integer PSD matrix can be represented as the sum of rank-$1$ matrices and one sporadic matrix, which is unimodularly equivalent to a matrix in the finite set $R$. Regarding prior work that inspired us, we mention [16] that contains a similar rank-1 decomposition structure for PSD $\{0,1\}$ matrices: a PSD $\{0,1\}$ matrix $X\in\mathcal{S}^{n}_{+}(\mathbb{Z})\cap\{0,1\}^{n\times n}$ satisfies $X=\sum_{i\in K}\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}$ for $\mathbf{x}_{i}\in\{0,1\}^{n}$, where $K$ is a finite index set. Similarly, [18] extended the results to PSD $\{0,\pm 1\}$ matrices: a PSD $\{0,\pm 1\}$ matrix $X\in\mathcal{S}^{n}_{+}(\mathbb{Z})\cap\{0,\pm 1\}^{n\times n}$ satisfies $X=\sum_{i\in K}\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}$ for $\mathbf{x}_{i}\in\{0,\pm 1\}^{n}$, where $K$ is a finite index set. Our results extend to all integer positive semidefinite matrices. For the second-order cone, we extended the construction of the Barning-Hall tree in [8] for the primitive extreme points (or Pythagorean tuples) to classify the sporadic points.
While it might be tempting to believe that these results hint that all conical semigroups are $(R,G)$-finitely generated for some finite set $R$ and some subgroup $G\subseteq\mathrm{GL}(N,\mathbb{Z})$, we conjecture the contrary:
Conjecture 1
There exists a conical semigroup $S_{C}$ that is not $(R,G)$-finitely generated for any choice of $R$ and $G$.
What is the significance of these results beyond their connections to classical geometry of numbers, lattices, and number theory? (See e.g., [14]). We motivate our interest about conical semigroups with two applications in optimization. In what follows, we assume that our cone $C\subset\mathbb{R}^{N}$ is full-dimensional.
The first application regards the notion of Chvátal-Gomory cuts which is useful in the branch-and-cut methods for integer programming. How much of this can be extended to conic integer programming? Given a linear map $\mathcal{A}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{m}\to\mathbb{R}^{N}$ and $\mathbf{c}\in\mathbb{R}^{N}$, we define a linear conical inequality (LCI) system as
$\mathrm{LCI}_{C}(\mathbf{c},\mathcal{A})\mathrel{\mathop{\ordinarycolon}}=\{\mathbf{x}\in\mathbb{R}^{m}\mathrel{\mathop{\ordinarycolon}}\mathbf{c}-\mathcal{A}(\mathbf{x})\in C\}$
where $\mathbf{c}\in\mathbb{Z}^{N}$ and $\mathcal{A}(\mathbb{Z}^{m})\subseteq\mathbb{Z}^{N}$. When $C$ is the cone of positive semidefinite matrices in $\mathcal{S}^{n}(\mathbb{R})$, then $N=\binom{n+1}{2}$ and $\mathcal{A}(\mathbf{x})=\sum_{i=1}^{m}x_{i}A_{i}$ for some matrices $A_{1},\dots,A_{m}\in\mathcal{S}^{n}(\mathbb{Z})$. This is known as a linear matrix inequality and defines a spectrahedron. An important concept for LCI is called total dual integrality (TDI), which has been well-known for polyhedral cones $C$ [13, 12] and recently extended to spectrahedral cones [7, 17]. We use $C^{*}$ to denote the dual cone of $C$, $\mathcal{A}^{*}$ to denote the adjoint linear map of $\mathcal{A}$, and give a definition for general cones here.
Definition 3
An LCI system $\mathbf{c}-\mathcal{A}(\mathbf{x})\in C$ is totally dual integral, if for any $\mathbf{b}\in\mathbb{Z}^{m}$, the dual optimization problem
$\min\quad$ $y(\mathbf{c})\quad\mathrm{s.t.}\quad\mathcal{A}^{*}(y)=\mathbf{b},\ y\in C^{*},$
whenever feasible, has an integer optimal solution $y^{*}\in C^{*}\cap\mathbb{Z}^{N}$.
To approximate the convex hull of $Z\mathrel{\mathop{\ordinarycolon}}=\mathrm{LCI}_{C}(\mathbf{c},\mathcal{A})\cap\mathbb{Z}^{m}$, a commonly used approach (quite similar to its polyhedral version) is to add Chvátal-Gomory (CG) cuts, which are defined as follows [17]. If $\mathbf{u}\in\mathbb{Z}^{m}$ is an integral vector and $v\in\mathbb{R}$ a real number such that the linear inequality $\mathbf{u}^{\mathsf{T}}\mathbf{x}\leq v$ is valid for all $x\in\mathrm{LCI}_{C}(c,\mathcal{A})$, then the inequality $\mathbf{u}^{\mathsf{T}}\mathbf{x}\leq\lfloor v\rfloor$ is valid for all $\mathbf{x}\in Z$ and called a CG cut. There are possibly infinitely many CG cuts so we define the (elementary) CG closure as
$\mathrm{CG}\text{-}\mathrm{cl}(Z)\mathrel{\mathop{\ordinarycolon}}=\bigcap_{\begin{subarray}{c}(\mathbf{u},v)\in\mathbb{Z}^{m}\times\mathbb{R}\mathrel{\mathop{\ordinarycolon}}\\ S\subseteq\{\mathbf{x}\mathrel{\mathop{\ordinarycolon}}\mathbf{u}^{\mathsf{T}}\mathbf{x}\leq v\}\end{subarray}}\left\{\mathbf{x}\in\mathbb{R}^{m}\mathrel{\mathop{\ordinarycolon}}\mathbf{u}^{\mathsf{T}}\mathbf{x}\leq\lfloor v\rfloor\right\}.$ (1)
Now take any linear function $w\in C^{*}$ such that $w(\mathbb{Z}^{N})\subseteq\mathbb{Z}$. Then, a CG cut can be generated by
$w\circ\mathcal{A}(\mathbf{x})\leq\lfloor w(\mathbf{c})\rfloor,$
as, by definition, $w\circ\mathcal{A}(\mathbb{Z}^{m})\in\mathbb{Z}$. Conversely, if the conical semigroup $S_{C^{*}}\mathrel{\mathop{\ordinarycolon}}=C^{*}\cap\mathbb{Z}^{N}$ is $(R,G)$-finitely generated, then we can get all CG cuts through $R$ and $G$ for our TDI LCI system. This is one of the nice consequences of this property.
Theorem 1.3
Suppose $C\subset\mathbb{R}^{N}$ is a full-dimensional convex cone such that $S_{C^{*}}\mathrel{\mathop{\ordinarycolon}}=C^{*}\cap\mathbb{Z}^{N}$ is $(R,G)$-finitely generated, and $\mathrm{LCI}_{C}(\mathbf{c},\mathcal{A})$ is TDI. Then the CG closure for $Z\mathrel{\mathop{\ordinarycolon}}=\mathrm{LCI}_{C}(\mathbf{c},\mathcal{A})\cap\mathbb{Z}^{m}$ can be described by
$\mathrm{CG}\text{-}\mathrm{cl}(Z)=\left\{\mathbf{x}\in\mathbb{R}^{m}\mathrel{\mathop{\ordinarycolon}}(g\cdot r)^{\mathsf{T}}\mathcal{A}(\mathbf{x})\leq\lfloor(g\cdot r)^{\mathsf{T}}\mathbf{c}\rfloor,\quad\forall\,r\in R,g\in G\right\}.$
The final application has to do with classical notions of integer rank [10]. Just like the notion of (real) rank of a linear system allows us to bound the number of nonzero entries in a solution of a linear system, we want to know how many elements are needed to decompose any element of a conical semigroup as a linear combination of generators with nonnegative integer coefficients. Suppose that our conical semigroup $S_{C}=C\cap\mathbb{Z}^{N}$ has an integer generating set $B$. For any element $s\in S_{C}$, there exist integer generators $b_{1},\dots,b_{m}\in B$ and $\lambda_{1},\dots,\lambda_{m}\in\mathbb{Z}_{\geq 1}$ such that $s=\sum_{i=1}^{m}\lambda_{i}b_{i}$, for some $m\geq 1$. The minimum number $m$ needed in the sum is called the integer Carathéodory rank (ICR) of $s$, and the maximum number over all $s\in S_{C}$ is the ICR of the conical semigroup $S_{C}$ or the cone $C$. We show an upper bound on the ICR that depends only on the dimension $N$. The proof is almost identical to the popular polyhedral result in [10, 22] but we must use the extreme point characterization of semi-infinite linear optimization [9] to allow infinite generating sets.
Theorem 1.4
Let $C\subset\mathbb{R}^{N}$ be an arbitrary pointed convex cone and $S_{C}\mathrel{\mathop{\ordinarycolon}}=C\cap\mathbb{Z}^{N}$. Then $\mathrm{ICR}(S_{C})\leq 2N-2$.
2 The Positive Semidefinite (PSD) Cone
Let $\mathcal{S}^{n}(\mathbb{Z})$ (resp. $\mathcal{S}^{n}(\mathbb{R})$) denote the set of $n\times n$ symmetric matrices of integer (resp. real) entries. For a matrix $X\in\mathcal{S}^{n}(\mathbb{Z})$, we say that $X$ is PSD (denoted as $X\succeq 0$) if and only if it is so when regarded as a real matrix $X\in\mathcal{S}^{n}(\mathbb{R})$. We denote $\mathcal{S}^{n}_{+}(\mathbb{Z})$ as the set of integer PSD matrices.
The group $\mathrm{GL}(n,\mathbb{Z})$ embeds into $\mathrm{GL}(N,\mathbb{Z})$ as follows. Given a matrix $U\in\mathrm{GL}(n,\mathbb{Z})$ and any $X\in\mathcal{S}^{n}(\mathbb{Z})$, we define the action $U\cdot X\mathrel{\mathop{\ordinarycolon}}=UXU^{\mathsf{T}}$. This action is a linear map and takes integer points in $\mathbb{Z}^{N}$ to integer points, and thus can be represented by the multiplication with a matrix in $\mathrm{GL}(N,\mathbb{Z})$. It is well-known that this group $\mathrm{GL}(n,\mathbb{Z})$ is finitely generated [23]. For the convenience of discussion, we still use the matrix $U\in\mathrm{GL}(n,\mathbb{Z})$ to denote this matrix multiplication in the subgroup of $\mathrm{GL}(N,\mathbb{Z})$.
2.1 Lemmas for $n\leq 5$ and $n=6$
The following integer rank-1 decomposition for PSD integer matrices is studied in [19]. We recast their arguments with a modern geometric perspective, and use it to extend the notion of $(R,G)$-finite generation to the PSD cone.
Lemma 1
If $n\leq 5$, then for any $X\in\mathcal{S}_{+}^{n}(\mathbb{Z})$, we can find a finite index set $K$ and vectors $\mathbf{x}_{i}\in\mathbb{Z}^{n}$, $i\in K$ such that
$X=\sum_{i\in K}\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}.$ (2)
To restate Definition 2 in the PSD case, we say an integer matrix $X\in\mathcal{S}^{n}(\mathbb{Z})$ is sporadic if there does not exist $\mathbf{x}\in\mathbb{Z}^{n}\setminus\{\mathbf{0}\}$ such that $X-\mathbf{x}\mathbf{x}^{\mathsf{T}}\succeq 0$. Lemma 1 is equivalent to the fact that there is no sporadic point in $\mathcal{S}_{+}^{n}(\mathbb{Z})$ when $n\leq 5$.
Proposition 1
There is no sporadic point in $\mathcal{S}_{+}^{n}(\mathbb{Z})$ if and only if every positive semidefinite integer matrix in $\mathcal{S}_{+}^{n}(\mathbb{Z})$ has an integer rank-1 decomposition.
For any matrix $X\in\mathcal{S}^{n}(\mathbb{R})$, we can define a convex set $C(X)\mathrel{\mathop{\ordinarycolon}}=\{\mathbf{x}\in\mathbb{R}^{n}\mathrel{\mathop{\ordinarycolon}}X-\mathbf{x}\mathbf{x}^{\mathsf{T}}\succeq 0\}$. Since $X-\mathbf{x}\mathbf{x}^{\mathsf{T}}\succeq 0$ if and only if, for any $\mathbf{v}\in\mathbb{R}^{n}$, $|\mathbf{v}^{\mathsf{T}}\mathbf{x}|^{2}\leq\mathbf{v}^{\mathsf{T}}X\mathbf{v}$, we see that $C(X)$ is a compact convex set that is symmetric about the origin. This provides another equivalent formulation of the integer rank-1 decomposition.
Proposition 2
For $X\in\mathcal{S}_{+}^{n}(\mathbb{Z})$, $X$ is sporadic if and only if $C(X)\cap\mathbb{Z}^{n}=\{\mathbf{0}\}$.
This provides a geometric perspective to our problem. Note that the set $C(X)$ is a (possibly degenerate) ellipsoid because
$X\succeq\mathbf{x}\mathbf{x}^{\mathsf{T}}\iff\begin{bmatrix}1&\mathbf{x}^{\mathsf{T}}\\ \mathbf{x}&X\end{bmatrix}\succeq 0\iff\mathbf{x}^{\mathsf{T}}X^{\dagger}\mathbf{x}\leq 1,(I-XX^{\dagger})\mathbf{x}=0,$
by the positive semidefiniteness of Schur complements, where $X^{\dagger}$ denotes the pseudoinverse of $X$. In the case where $\det(X)>0$,
$C(X)=\{\mathbf{x}\in\mathbb{R}^{n}\mathrel{\mathop{\ordinarycolon}}\mathbf{x}^{\mathsf{T}}\mathrm{adj}(X)\mathbf{x}\leq\det(X)\}\text{ with }\mathrm{vol}(C(X))=V_{n}\sqrt{\det(X)}$
where $\mathrm{adj}(X)$ is the adjugate of $X$ satisfying $\mathrm{adj}(X)=\det(X)X^{-1}$ and $V_{n}\mathrel{\mathop{\ordinarycolon}}=\pi^{n/2}/\Gamma(\frac{n}{2}+1)$ is the volume of the unit $n$-ball. Note that $C(X)$ is not necessarily full-dimensional if $X$ is rank deficient. The degenerate case for rank 1 is characterized by the following proposition.
Proposition 3
Suppose $X\in\mathcal{S}_{+}^{n}(\mathbb{Z})$, and $\mathrm{rank}(X)=1$, then $X=\lambda\mathbf{x}\mathbf{x}^{\mathsf{T}}$, where $\mathbf{x}\in\mathbb{Z}^{n}$ and $\lambda\in\mathbb{Z}_{\geq 1}$.
From Proposition 3, we can directly prove the case for $n=2$ using Minkowski’s first Theorem (for example, see [11]): if $\mathrm{vol}(C(X))>4$, then there is a nonzero integral point.
Proposition 4
Lemma 1 holds for $n=2$.
In the general degenerate case, we can reduce the problem to one involving full-rank matrices of some lower dimension.
Lemma 2
Let $X\in\mathcal{S}^{n}(\mathbb{Z})$. If $r=\mathrm{rank}(X)<n$, then $X$ is unimodularly equivalent to
$\begin{bmatrix}\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\hat{X}\end{bmatrix},$
for some $\hat{X}\in\mathbb{Z}^{r\times r}$, $\mathrm{rank}(\hat{X})=r$.
To prove Lemma 1, we need to use a more sophisticated method based on the Hermite constant [21]
$\gamma_{n}\mathrel{\mathop{\ordinarycolon}}=\left(\max_{A\succ 0}\frac{\lambda_{1}(A)}{(\det(A))^{\frac{1}{n}}}\right)^{n},\text{ where }\lambda_{1}(A)=\min_{\mathbf{x}\in\mathbb{Z}^{n}\setminus\{\bf 0\}}(\mathbf{x}^{\mathsf{T}}A\mathbf{x}).$
Remark 1
Hermite gives a bound $\gamma_{n}\leq(\frac{4}{3})^{\frac{n(n-1)}{2}}$. The exact value of $\gamma_{n}$ is only known for $n\leq 8$ and $n=24$.
$n$
2
3
4
5
6
7
8
24
$\gamma_{n}$
$\frac{4}{3}$
2
2
8
$\frac{64}{3}$
64
256
$4^{24}$
Proof (for Lemma 1)
The case $n=1$ follows from Proposition 3. We will show that $C(X)\cap\mathbb{Z}^{n}\neq\{\mathbf{0}\}$ for $2\leq n\leq 5$, where $C(X)=\{\mathbf{x}\in\mathbb{R}^{n}\mathrel{\mathop{\ordinarycolon}}\mathbf{x}^{\mathsf{T}}\mathrm{adj}(X)\mathbf{x}\leq\det(X)\}$. By the definition of the Hermite constant, we have
$\min_{\mathbf{x}\in\mathbb{Z}^{n}\setminus\{\bf 0\}}(\mathbf{x}^{\mathsf{T}}\mathrm{adj}(X)\mathbf{x})\leq(\gamma_{n}\det(\mathrm{adj}(X)))^{\frac{1}{n}}=(\gamma_{n}(\det(X))^{n-1})^{\frac{1}{n}}.$
For $n=2,3,4,5$, we have $\frac{n^{n}}{(n-1)^{n-1}}>\gamma_{n}$. By taking the derivative with respect to $\det(X)$, we know that $\frac{(\det(X)+1)^{n}}{(\det(X))^{n-1}}\geq\frac{n^{n}}{(n-1)^{n-1}}$. Thus, $\gamma_{n}<\frac{(\det(X)+1)^{n}}{\det(X)^{n-1}}$. Therefore,
$\min_{\mathbf{x}\in\mathbb{Z}^{n}\setminus\{\bf 0\}}(\mathbf{x}^{\mathsf{T}}\mathrm{adj}(X)\mathbf{x})\leq(\gamma_{n}\det(\mathrm{adj}(X)))^{\frac{1}{n}}=(\gamma_{n}\det(X)^{n-1})^{\frac{1}{n}}<\det(X)+1.$
Because $x^{\mathsf{T}}\mathrm{adj}(X)x,\det(X)\in\mathbb{Z}$ for any $x\in\mathbb{Z}^{n}$, we have
$\min_{\mathbf{x}\in\mathbb{Z}^{n}\setminus\{\bf 0\}}(\mathbf{x}^{\mathsf{T}}\mathrm{adj}(X)\mathbf{x})\leq\det(X),$
which implies that $C(X)\cap(\mathbb{Z}^{n}\setminus\{\mathbf{0}\})\neq\emptyset$. Lemma 1 now follows from Propositions 2 and 1, and Lemma 2. ∎
The argument used to prove Lemma 1 fails for $n\geq 6$, but it implies that the determinant of the sporadic matrices is bounded by a constant only dependant on $n$. For example, in the case of $n=6$, the argument only fails when $3\leq\det(X)\leq 14$; for $n=7$, it only fails when $2\leq\det(X)\leq 56$, and for $n=8$, it only fails when $1\leq\det(X)\leq 247$. We summarize this observation in the following corollary.
Corollary 1
If $X\in\mathcal{S}_{+}^{n}(\mathbb{Z})$ is sporadic, then $\det(X)<\gamma_{n}$.
A sporadic matrix for $n=6$ was initially found in [19].
Proposition 5
In $n=6$, the matrix $M$ is sporadic, i.e., $C(M)\cap\mathbb{Z}^{n}=\{\mathbf{0}\}$.
[…]
Moreover, in [15], it is shown that for $n=6$, $M$ is the unique sporadic matrix under unimodular equivalence. Using this fact, we have the following Lemma.
Lemma 3
If $n=6$, then for any $X\in\mathcal{S}_{+}^{n}(\mathbb{Z})$,
$X=\sum_{i\in K}\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}+Y$
for $\mathbf{x}_{i}\in\mathbb{Z}^{n}$ and $Y$ unimodularly equivalent to $M$, where $K$ is a finite index set.
2.2 Proof of Theorem 1.1
Proof
We know that the primitive extreme points are generated from the group $\mathrm{GL}(n,\mathbb{Z})$ that acts on $\{\mathbf{e}_{1}\mathbf{e}_{1}^{\mathsf{T}}\}$. The finiteness of the index set $K$ follows from similar argument in Proposition 1. We only need to prove that the sporadic points are generated from the group $\mathrm{GL}(n,\mathbb{Z})$ on a finite set $R$.
Corollary 1 shows that for any sporadic matrix $X$, $\det(X)<\gamma_{n}$. By [21, Theorem 2.4], there exists a constant $\alpha_{n}>0$ depending only on $n$, such that for any positive definite matrix $X\in\mathcal{S}^{n}(\mathbb{Z})$, there is a unimodularly equivalent matrix $X^{\prime}$ of $X$ with diagonal entries satisfy
$\prod_{i=1}^{n}X^{\prime}_{ii}\leq\alpha_{n}\det(X^{\prime})=\alpha_{n}\det(X)<\alpha_{n}\gamma_{n}.$
Because $X^{\prime}\in\mathcal{S}^{n}(\mathbb{Z})$ is positive definite, $X^{\prime}_{ii}\geq 1$ and thus is bounded from above. From this we see that there are only finitely many possibilities for such $X^{\prime}$ because each off-diagonal entry must satisfy $|X^{\prime}_{ij}|^{2}\leq X^{\prime}_{ii}X^{\prime}_{jj}$ for any $1\leq i,j\leq n$.
The two special cases $n\leq 5$ and $n=6$ follow from Lemma 1 and 3. ∎
3 The Second-Order Cone (SOC)
In this section, we will let $T_{n}$ be the conical semigroup $\text{SOC}(n)\cap\mathbb{Z}^{n}$ where
${\small\text{SOC}(n)\mathrel{\mathop{\ordinarycolon}}=\left\{\mathbf{x}\in\mathbb{R}^{n}\mathrel{\mathop{\ordinarycolon}}0\leq\sqrt{x_{1}^{2}+\dots+x_{n-1}^{2}}\leq x_{n}\right\}.}$
Additionally, for $\mathbf{a},\mathbf{b}\in\mathbb{R}^{n}$, consider the bilinear form
$\langle\mathbf{a},\mathbf{b}\rangle\mathrel{\mathop{\ordinarycolon}}=a_{1}b_{1}+a_{2}b_{2}+\dots+a_{n-1}b_{n-1}-a_{n}b_{n}.$
In this quadratic space, the reflection in vector $\mathbf{w}$ is defined as $\mathbf{x}\rightarrow\mathbf{x}-2\frac{\langle\mathbf{x},\mathbf{w}\rangle}{\langle\mathbf{w},\mathbf{w}\rangle}\mathbf{w}$.
Definition 4
Let $P_{i,j}$ be the permutation matrix that swaps the $i$-th and $j$-th columns and define $Q_{k}$ be the matrix determined by
$(Q_{k})_{i,j}=\begin{cases}-1&\text{ if }i=j=k\\ 1&\text{ if }i=j\neq k\\ 0&\text{ if }i\neq j.\end{cases}$
For $n=3$, let $A_{3}$ denote the matrix associated with the reflection in the vector $(1,1,1)$. For $4\leq n\leq 10$, let $A_{n}$ denote the matrix associated with the reflection in the vector $(1,1,1,0,\dots,0,1)$ also associated to this bilinear form:
$A_{3}=\begin{pmatrix}-1&-2&2\\ -2&-1&2\\ -2&-2&3\\ \end{pmatrix},\hskip 8.53581ptA_{n}=\begin{pmatrix}\begin{matrix}0&-1&-1\\ -1&0&-1\\ -1&-1&0\end{matrix}&\vline&\mbox{\Large 0}&\vline&\begin{matrix}1\\ 1\\ 1\end{matrix}\\ \hline\cr\mbox{\Large 0}&\vline&\text{I}_{n-4}&\vline&\mbox{\Large 0}\\ \hline\cr\begin{matrix}-1&-1&-1\end{matrix}&\vline&\mbox{\Large 0}&\vline&2\end{pmatrix}$
We define the matrix $A_{n}^{+}=Q_{1}Q_{2}\dots Q_{n-1}A_{n}$. Note that $A_{n}$ is unimodular.
Elements $\mathbf{s}\in T_{n}$ such that $\langle\mathbf{s},\mathbf{s}\rangle=\bf 0$ belong to the boundary of $T_{n}$, and we will denote the set of these points as $\partial T_{n}$. In number theory, these points are called Pythagorean tuples. In [8], they proved that the set of primitive Pythagorean tuples, denoted as $\operatorname{ext}^{p}(T_{n})$, is generated by finitely many matrices acting on a finite set $R$ for $3\leq n\leq 10$.
Lemma 4 (Theorem 1 in [8])
For $3\leq n\leq 10$, $\operatorname{ext}^{p}(T_{n})=\cup_{r\in R}G\cdot r$, where the group
$G=\left\langle A_{n},Q_{1},\dots,Q_{n-1},P_{1,2},P_{1,3},\dots P_{1,n-1}\right\rangle$
and the sets
  1. 1.
    $R=\left\{(1,0,\dots,0,1)^{\mathsf{T}}\right\}$ for $103\leq n<$ 10,
  2. 2.
    $R=\left\{(1,0,0,0,0,0,0,0,0,1)^{\mathsf{T}},(1,1,1,1,1,1,1,1,1,3)^{\mathsf{T}}\right\}$ for $n=10$,
where $G$ acts on $R$ by left multiplication.
Note that $G$ maps $\partial T_{n}$ to $\partial T_{n}$.
We will begin this section by discussing the structural properties of the sporadic points of $T_{n}$. Then we use the structural properties of Pythagorean tuples and sporadic points to prove Theorem 1.2.
3.1 Sporadic Points of $\text{SOC}(n)\cap\mathbb{Z}^{n}$
In this section, we will begin by restating the definition of sporadic points in the case of $\text{SOC}(n)$ and offer two partial characterizations of sporadic elements of $\text{SOC}(n)$.
Definition 5
We call a point $\mathbf{s}\in T_{n}$ sporadic if there is no integral point $\mathbf{p}$ such that $\langle\mathbf{p},\mathbf{p}\rangle=0$ and $\mathbf{s}-\mathbf{p}\in T_{n}$.
Just as the group $G$ takes elements of $\partial T_{n}$ to $\partial T_{n}$, the group $G$ will take sporadic elements to sporadic elements. This closure ensures that our action by $G$ on the semigroup $T_{n}$ is well-defined.
Lemma 5
Let $\mathbf{s}\in T_{n}$.
  1. 1.
    Then, $A_{n}^{+}\mathbf{s}$ and $(A_{n}^{+})^{-1}\mathbf{s}$ are both in $T_{n}$.
  2. 2.
    If $\mathbf{s}$ is sporadic, then $A_{n}^{+}\mathbf{s}$ and $(A_{n}^{+})^{-1}\mathbf{s}$ are both sporadic.
Next, we will provide some Lemmas about the properties of sporadic points necessary to prove Theorem 1.2. For a detailed exposition of the technical proofs, please refer to the extended version. Lemmas 6 and 7 show that sporadic points are close to the boundary, $\partial T_{n}$.
Lemma 6
Suppose $\mathbf{s}\in T_{n}$ is a primitive sporadic with nonnegative entries such that $s_{n}>1$ and $s_{i}\neq 0$ for some $i\in[n-1]$. Then,
$s_{n}=\left\lceil\sqrt{s_{1}^{2}+s_{2}^{2}+\dots+s_{n-1}^{2}}\right\rceil$
Lemma 7
Let $\mathbf{s}\in T_{n}$. If $\langle\mathbf{s},\mathbf{s}\rangle=-1$, then $\mathbf{s}$ is sporadic. When $73\leq n<$ 7, we also have the converse: if $\mathbf{s}$ is sporadic, then $\langle\mathbf{s},\mathbf{s}\rangle=-1$.
Inspired by the structure of Pythagorean tuples, we analyze the set of sporadic points that remain at the same height in $T_{n}$ after multiplication by $(A_{n}^{+})^{-1}$. Let $(p)_{n}$ denotes the $n^{\text{th}}$ coordinate of $p$,
Lemma 8
Let $n\leq 10$. Suppose $\mathbf{s}\in T_{n}$ is a primitive sporadic such that $s_{1}\geq\dots\geq s_{n-1}\geq 0$ and $s_{n}>1$. The following list of tuples are the only such $\mathbf{s}$ where $((A_{n}^{+})^{-1}\mathbf{s})_{n}=s_{n}$.
  • For $n=7$, we have the following tuple: $(1,1,1,1,1,1,3)$.
  • For $n=8$, we have the following tuples: $(1,1,1,1,1,1,1,3)$, $(1,1,1,1,1,1,0,3)$.
  • For $n=9$, we have the following tuples:
    $(1,1,1,1,1,1,1,1,3),(1,1,1,1,1,1,1,0,3),(1,1,1,1,1,1,0,0,3),(2,2,2,2,2,2,2,1,6).$ .
  • For $n=10$, we have the following tuples:
    $(1,1,1,1,1,1,1,1,0,3),$ $\hskip 8.53581pt(1,1,1,1,1,1,1,0,0,3),$ $\hskip 8.53581pt(1,1,1,1,1,1,0,0,0,3),$
    $(2,2,2,2,2,2,2,2,1,6),$ $\hskip 8.53581pt(2,2,2,2,2,2,2,1,0,6).$
Then we show that besides the points listed in Lemma 8, every other sporadic points will reduce to a strictly lower height after multiplication by $(A_{n}^{+})^{-1}$.
Lemma 9
Let $\mathbf{s}\in T_{n}$ be sporadic with nonnegative entries such that $s_{1}\geq s_{2}\geq\dots\geq s_{n-1}$, $s_{1}\geq 1$ and $3\leq n\leq 10$. For $s$ not listed in Lemma 8,
$((A_{n}^{+})^{-1}\mathbf{s})_{n}<(\mathbf{s})_{n}.$
Proof
This is equivalent to showing that $-s_{1}-s_{2}-s_{3}+2s_{n}<s_{n}$, or rather $s_{n}<s_{1}+s_{2}+s_{3}.$ The case of $n=3$ reduces to the inequality $s_{3}<s_{1}+s_{2}$. By Lemma 6, we have that
$s_{n}$ $=\bigg{\lceil}\sqrt{s_{1}^{2}+s_{2}^{2}+\dots+s_{n-1}^{2}}\bigg{\rceil}$
$\leq\bigg{\lceil}\sqrt{s_{1}^{2}+s_{2}^{2}+s_{3}^{2}+2s_{1}s_{2}+2s_{1}s_{3}+2s_{2}s_{3}}\bigg{\rceil}$ (3)
$=\bigg{\lceil}\sqrt{(s_{1}+s_{2}+s_{3})^{2}}\bigg{\rceil}=s_{1}+s_{2}+s_{3},$
where the inequality follows from the order $s_{1}s_{2}\geq s_{1}s_{3}\geq s_{2}s_{3}\geq s_{4}^{2}\geq...\geq s_{n-1}^{2}$ and $n\leq 10$. As $s$ is not one of the tuples listed in Lemma 8, the inequality (3) can be made strict. Therefore, $s_{n}<s_{1}+s_{2}+s_{3}$, which implies that $(A_{n}^{+})^{-1}\mathbf{s}$ sits at a strictly lower height in the cone than $\mathbf{s}$. ∎
3.2 Proof of Theorem 1.2
We now present a complete formulation of Theorem 1.2 followed by its proof.
Theorem 3.1
For dimension $3\leq n\leq 10$, the conical semigroup $\mathrm{SOC}(n)\cap\mathbb{Z}^{n}$ is $(R,G)$-finitely generated by
$G=\left\langle A_{n}^{+},Q_{1},\dots,Q_{n-1},P_{1,2},P_{1,3},\dots P_{1,n-1}\right\rangle$
and a finite set $R$. More specifically,
  1. 1.
    If $3\leq n\leq 6$, then $R=\left\{(1,0,\dots,0,1)^{\mathsf{T}},(0,\dots,0,1)^{\mathsf{T}}\right\}$.
  2. 2.
    If $n=7$, then
    $R=\left\{(1,0,0,0,0,0,1)^{\mathsf{T}},(0,0,0,0,0,0,1)^{\mathsf{T}},(1,1,1,1,1,1,3)^{\mathsf{T}}\right\}.$
  3. 3.
    If $n=8$, then
    $R=\Big{\{}(1,0,0,0,0,0,0,1)^{\mathsf{T}},(0,0,0,0,0,0,0,1)^{\mathsf{T}},(1,1,1,1,1,1,1,3)^{\mathsf{T}},(1,1,1,1,1,1,0,3)^{\mathsf{T}}\Big{\}}.$
  4. 4.
    If $n=9$, then
[…]
    $(1,1,1,1,1,1,1,0,3)^{\mathsf{T}},(1,1,1,1,1,1,0,0,3)^{\mathsf{T}},(2,2,2,2,2,2,2,1,6)^{\mathsf{T}}\Big{\}}.$
  5. 5.
    If $n=10$, then
    $R=\Big{\{}(1,0,0,0,0,0,0,0,0,1)^{\mathsf{T}},(1,1,1,1,1,1,1,1,1,3)^{\mathsf{T}},(0,0,0,0,0,0,0,0,0,1)^{\mathsf{T}},$
    $(1,1,1,1,1,1,1,1,0,3)^{\mathsf{T}},(1,1,1,1,1,1,1,0,0,3)^{\mathsf{T}},(1,1,1,1,1,1,0,0,0,3)^{\mathsf{T}},$
    $(2,2,2,2,2,2,2,2,1,6)^{\mathsf{T}},(2,2,2,2,2,2,2,1,0,6)^{\mathsf{T}}\Big{\}}.$
Proof
This follows directly from Lemma 9 and that fact that $(0,0,\dots,0,1)$ is the sporadic of minimal height in this cone. Let $\mathbf{s}\in T_{n}$. If $\mathbf{s}$ is not sporadic, we can represent it as
$\mathbf{s}=\lambda_{1}\mathbf{p}_{1}+\lambda_{2}\mathbf{p}_{2}+\dots+\lambda_{k}\mathbf{p}_{k}+\lambda\mathbf{p}$ (4)
where $\lambda,\lambda_{i}\in\mathbb{Z}_{\geq 0}$, each $\mathbf{p}_{i}$ is a primitive Pythagorean tuple and $\mathbf{p}$ is sporadic. By Lemma 4, each $\mathbf{p}_{i}$ can be decomposed as $\mathbf{p}_{i}=G_{i}(1,0,\dots,0,1)^{\mathsf{T}}$ when $103\leq n<$ 10 or $\mathbf{p}_{i}=G_{i}(1,0,\dots,0,1)^{\mathsf{T}}+\tilde{G_{i}}(1,1,\dots,1,3)^{\mathsf{T}}$ when $n=10$ where each $G_{i},\tilde{G_{i}}\in G$. It remains to consider the sporadic $\mathbf{p}$. Given any primitive sporadic tuple $\mathbf{s}$, we can recover an element of $R$ as follows:
  1. 1.
    Multiply $\mathbf{p}$ by the appropriate permutation matrices $P_{i,j}$ and sign changing matrices $Q_{j}$ so that $\mathbf{p}$ has nonnegative entries and $p_{1}\geq\dots\geq p_{n-1}$. Call this resulting vector $\mathbf{p}^{\prime}$.
  2. 2.
    Multiply $\mathbf{p}^{\prime}$ by $(A_{n}^{+})^{-1}$ and repeat step 1 as necessary. By Lemma 9, the height of the resulting vector will be strictly lower then that of the vector we started with or the resulting vector will belong to $R$.
  3. 3.
    Repeat step 2 until the resulting vector $\mathbf{r}$ belongs to $R$. By Lemma 8, the only possibilities for the resulting vector belong to $R$.
This process gives the equality $\mathbf{r}=G_{1}\dots G_{k}\mathbf{p}$. If we let $G^{\prime}=G_{1}\dots G_{k}$, then we have $(G^{\prime})^{-1}\mathbf{r}=\mathbf{p}$. Therefore, for $3\leq n\leq 10$, the conical semigroup $\text{SOC}(n)$ is $(R,G)$-finitely generated by the claimed $R$ and $G$. ∎
When $n=9$, the primitive sporadic point $(2,2,2,2,2,2,2,1,6)$ can be written as the sum of two sporadic points with smaller heights:
$(2,2,2,2,2,2,2,1,6)=(1,1,1,1,1,1,1,0,3)+(1,1,1,1,1,1,0,0,3).$
We can similarly decompose $(2,2,2,2,2,2,2,2,1,6)$ and $(2,2,2,2,2,2,2,1,0,6)$ for $n=10$. In this sense, these sporadic points fail to be minimal. Thus, if we remove them from the set of roots $R$, our semigroup $S$ remains $(R,G)$-finitely generated. However, when we remove these point from our root sets, our decomposition in equality (4) requires modification and we must allow for multiple sporadic points in the expression.
Remark 2
Lastly, it is worth noting that inequality (3) would fail in dimensions larger than 10. Thus, this line of argumentation would fail to produce results for $n>10$.